Consider a schedule with 4 transactions with 1, 2, 3, and 4 operations respectively. Calculate the possible

number of

a) Serial schedules

b) Non serial schedules


1
Expert's answer
2021-11-11T17:35:05-0500

Total Number of Schedules,

= (1 + 2 + 3 + 4)! / (1! + 2! + 3! + 4!) = 10! / 33 = 109963

Number of Serial Schedules,
= 4! = 24

Number of Non-Serial Schedules,
= Total Number of Schedules - Number of Serial Schedules 
= 109963 - 24 
= 109939